OpenStudy (trojanpoem):

(1) Find binomial expansion for: (2) Find the expansion for the given function using Maclaurin series for common functions: sqrt( 1+ sinx)

OpenStudy (trojanpoem):

\[\frac{ x^2 + 16x + 2 }{ x^3 + 23x^2 + 20x - 4}\]

OpenStudy (anonymous):

you need to factor it

OpenStudy (trojanpoem):

@SkepticGod , the roots aren't integers.

OpenStudy (anonymous):

I think I remeber it now so isn't it 1+sin^2=cos^2x ?

OpenStudy (anonymous):

I think the first thing you have to do is solve for x and then plug them in for the "x"

OpenStudy (trojanpoem):

sqrt(1+sinx) sinx = sqrt(1+cos^2x) sqrt(1+ sqrt(1+cos^2x)) = useless, If I understood you well.

OpenStudy (trojanpoem):

We need to get rid of the root. I tried using the common expansion for (1+x)^m than plugging x = sinx Got it, but long solution.

OpenStudy (anonymous):

sorry nevermind lol

OpenStudy (anonymous):

I mean there's nothing I could really expand further because both denominator and nominator are prime

OpenStudy (trojanpoem):

@SkepticGod , thanks at least you're thinking with me.

OpenStudy (anonymous):

no problem! :D seem like I knew it but I haven't learn that Taylor/Maclaurin polynomial and series so..

OpenStudy (trojanpoem):

@SkepticGod , did you learn binomial ?

OpenStudy (anonymous):

yea i did when I was in High school

OpenStudy (trojanpoem):

@SkepticGod , Do you know that when the power is negative , it's infinite series ?

OpenStudy (anonymous):

what i know is expanding ding binomial thats all, wish i could help but too bad i haven't expert on series

OpenStudy (mathmale):

Trojo: Have you done similar problems requiring binomial expansions? At first glance, this one is a bear.

OpenStudy (trojanpoem):

@mathmale , I 've been doing binomial expansions for about 1 year ( + ve real number power) and for 0.5 year in expansions with negative , rational numbers.

OpenStudy (mathmale):

Have you tried every approach you can think of to factor numerator and denominator?

OpenStudy (trojanpoem):

@mathmale , Tried grouping (failed), finding roots using calc ( none integers) long division doesn't give any beneficial results.

OpenStudy (trojanpoem):

@mathmale , I even thought of partial fractions xD

OpenStudy (mathmale):

Have you tried dividing the numerator into the denominator?

OpenStudy (trojanpoem):

@mathmale I got 1- something by adding , subtracting. which was of no use. divinding ? you will get reminder in the denominator.

OpenStudy (mathmale):

Do you have your calculus book with y ou? I'll go get mine. Need a quick review of the binom. thm.

OpenStudy (mathmale):

Excuse me, binomial expansions. Sorry.

OpenStudy (anonymous):

Binomial Theorem: just in case (x+y)^n = \[\sum_{k=0}^{n}\left(\begin{matrix}n \\ k\end{matrix}\right) x^{n-k} y ^{k}\] where, \[\left(\begin{matrix}n \\ k\end{matrix}\right)=\frac{ n! }{ k!(n-k)! }\]

OpenStudy (trojanpoem):

@SkepticGod , Your formula won't work with a negative power like in our case.

OpenStudy (trojanpoem):

Replace n with -n

OpenStudy (mathmale):

thanks, Skeptic. Just realized something. Binomial expansions involve expanding powers of binomials, such as in \[(1+x)^m\]

OpenStudy (mathmale):

The complicated rational function you're working on certainly does not look like a binomil at first glance. Am I missing something?

Parth (parthkohli):

it's an infinite series... not a binomial expansion. he meant that.

OpenStudy (trojanpoem):

Totally, correct. We have to do some magic and turn it into good thing.

Parth (parthkohli):

the binomial theorem for any index...

OpenStudy (mathmale):

(1) Find binomial expansion for: (2) Find the expansion for the given function using Maclaurin series for common functions: sqrt( 1+ sinx) Trojan: what say you?

OpenStudy (mathmale):

Where does your "sqrt(1+sin x) come from?

Parth (parthkohli):

\[(1+y)^n = 1 + ny + \frac{n(n+1)}{2}y^2 + \cdots \]

OpenStudy (trojanpoem):

I don't know mate. That's why I am asking. the second one we have to get rid of the square root or use this method: using the common expansion for (1+x)^m than plugging x = sinx Got it, but long solution.

OpenStudy (trojanpoem):

The sqrt(1+sinx) is another question. If one is solved and I will take the other and post a new question.

OpenStudy (trojanpoem):

Did you notice that the question had been here for 4 hours ?

OpenStudy (mathmale):

Trojan, do you agree or disagree with Parth's comment, in which he says you meant an "infinite series" (which could be either a Taylor or Maclaurin series?

OpenStudy (trojanpoem):

@mathmale , I totally agree with parthkohli , he understand what I mean. The binomial is infinite as the power is negative (x^3+..)^-1

OpenStudy (trojanpoem):

See this: http://mathworld.wolfram.com/NegativeBinomialSeries.html

OpenStudy (mathmale):

But does the given denominator have the format (a+b)?

OpenStudy (trojanpoem):

Nope.

OpenStudy (trojanpoem):

Maybe if you can find the roots as 1+ sqrt(3)/3 or something I can play around with partial fractions.

OpenStudy (mathmale):

\[(a+b)^3=a^3+3a^2b+3ab^2 + b^3.\] and I don't see that your denom has that form. Thus, as interesting as that Wolfram excerpt is, it may not aply hear.

OpenStudy (mathmale):

apply here, I meant.

OpenStudy (mathmale):

Partial fractions would work, probably, if and only if y ou can factor the denominator.

OpenStudy (mathmale):

I've begun trying to factor it using synth. div., but have not finished. Have you tried factoring by this method?

OpenStudy (trojanpoem):

I may use the formula if I want .. ((x^3 + 23x^2) + (20x - 4))^-1 Let x^3 + 23x^2 be z Let 20x -4 be y (z + y)^-1 use the formula replace and keep expanding each one considering an infinite series I will die before archiving my aim.

OpenStudy (mathmale):

Actually, Trojo, you need only write out the first three or so terms to illustrate the pattern. I'd much rather that you didn't die.

OpenStudy (mathmale):

Again: Have you tried factoring the den. using synth. div? I've tried the possible roots 1, -1, 2 and -2, and have found that none lead to factors of the den.

OpenStudy (mathmale):

Could we temporarily abandon binom. expansions and focus on Maclaurin expansins? Have you tried using the def'n of the Mac expansion to come up with the first few terms of the Mac expansion of the given rational function?

OpenStudy (trojanpoem):

Imagine : (x^2 + 16x + 2) ( z+y)^-1 expand (z+y)^-1: (z+y)^-1 = 1/z * (1 + y/z)^-1 = 1/z * [ 1 + y/z + (y/z)^2 + (y/z)^3 + .......] = [1/z + y/z^2 + y^2/z^3 +y^3/z^4 + ....] [1/z + y/z^2 + y^2/z^3 +y^3/z^4 + ....](x^2 + 16x + 2) [ 1/z x^2 + x^2 * y/z^2 + y^2/z^3 * x^2 + .... ] +[ 16x * 1/z * 16 x * y/z^2 + ....] + [ 2 * 1/z + y/z^2 * 2 + 2 * y^2/z^3 + ...] Then replace z, y with actual values and simplify

OpenStudy (trojanpoem):

@mathmale , Hahaha I used the fking damn calculator and got this: root 1 : -22.08625958 root 2: 0.1675002457 root 3: -1.0871240664

OpenStudy (mathmale):

I see what you've done at the beginning: (x^2 + 16x + 2) ( z+y)^-1 but lose you as you expand this idea further. So you are saying that your den has 3 real roots? If that's the case, you can convert each into a factor and then use Partial Fractions.

OpenStudy (mathmale):

But what a messy bit o' work this will turn into!

ganeshie8 (ganeshie8):

\[\frac{ x^2 + 16x + 2 }{ x^3 + 23x^2 + 20x - 4} \\~\\=-\frac{1}{4} (x^2 + 16x + 2 ) (1-g)^{-1} \\~\\=-\frac{1}{4} (x^2 + 16x + 2 ) (1+g+g^2+\cdots )\] where \(g = \frac{1}{4}(x^3+23x^2+20x) \)

ganeshie8 (ganeshie8):

just write first few terms in the series and move to next problem :)

Parth (parthkohli):

is that an OK solution? i thought exactly the same thing but... looks too "unsimplified"?

ganeshie8 (ganeshie8):

I would write at least upto the x^4 term and put \(\cdots \)

OpenStudy (mathmale):

Quite an interesting substitution. Would you separate the given rational function into 3, all with the same denom? Then mult. the binomial expansion by x^2, by 16x and by 2, one by one?

OpenStudy (trojanpoem):

@ganeshie8 , How did you get rid of the 4 and got the 1/4. the question require simplification.

OpenStudy (trojanpoem):

@mathmale , Same as finding the expansion and multiplying times (x^2 +16x + 2)

OpenStudy (trojanpoem):

@ganeshie8 , true, i need the first 3 terms.

OpenStudy (mathmale):

Yes, that's familiar and makes sense.

OpenStudy (trojanpoem):

@ParthKohli , We need to simplify , I would 've solved it years ago.

ganeshie8 (ganeshie8):

\[\frac{ x^2 + 16x + 2 }{ x^3 + 23x^2 + 20x - 4} \\~\\ \frac{ x^2 + 16x + 2 }{-4[ 1 - \frac{1}{4}(x^3 + 23x^2 + 20x )]} \\~\\ =-\frac{1}{4} (x^2 + 16x + 2 ) (1-g)^{-1} \\~\\=-\frac{1}{4} (x^2 + 16x + 2 ) (1+g+g^2+\cdots )\] where \(g = \frac{1}{4}(x^3+23x^2+20x) \)

OpenStudy (mathmale):

So, approaching this problem that way, we do have a binomail to expand:\[(1-g)^{-1}\]

OpenStudy (trojanpoem):

@ganeshie8 , nice trick indeed. Now simplify please xD

OpenStudy (trojanpoem):

@ParthKohli , your turn. find the 3 terms. Alexander isn't here to multiply.

OpenStudy (trojanpoem):

@mathmale , yeah and a lot of inner binomials to expand too.

OpenStudy (anonymous):

I didn't understand this at all... D:

OpenStudy (mathmale):

Seems that we'd get a sum \[x^2(1-g)^{-1}+16x(1-g)^{-1}+2(1-g)^{-1}\]

Parth (parthkohli):

how would you find the degree zero term there?

Parth (parthkohli):

1, g, g^2, ... all generate degree zero terms.

OpenStudy (trojanpoem):

Hmm theoretically. maybe just 2

OpenStudy (trojanpoem):

@ParthKohli , remember that g= x^3 + 23x^2 + 20x which will give degree 1 +

OpenStudy (mathmale):

Parth: Same comment from me on that. g(0)=0; (1-g) is not zero.

Parth (parthkohli):

sorry :(

OpenStudy (trojanpoem):

@ParthKohli , Go multiply it's easy lol you will get expansion of (x^3 + 23x^2 + 20x)^2 < Ok ? ((x^3+23x^2) + 20x)^2 After expanding you will get (x^3 + 23x^2)^n (n differs) Now expand again and you've finished the first expansion go to the next.

ganeshie8 (ganeshie8):

why do you want to expand (x^3 + 23x^2 + 20x)^2

ganeshie8 (ganeshie8):

unless your prof is a psycho

Parth (parthkohli):

there has to be a simpler way

OpenStudy (trojanpoem):

@ganeshie8 to get power 20 x , Maybe he got a trick. By the way do you know simpler solution for the second ?

OpenStudy (trojanpoem):

@ParthKohli , Correct.

OpenStudy (mathmale):

Trojo: I'm still wondering whether you've looked thru your text, notes, etc., for concrete examples of how to expand rational functions such as yours into a binomial series.

ganeshie8 (ganeshie8):

\[\frac{ x^2 + 16x + 2 }{ x^3 + 23x^2 + 20x - 4} \\~\\ \frac{ x^2 + 16x + 2 }{-4[ 1 - \frac{1}{4}(x^3 + 23x^2 + 20x )]} \\~\\ =-\frac{1}{4} (x^2 + 16x + 2 ) (1-g)^{-1} \\~\\=-\frac{1}{4} (x^2 + 16x + 2 ) (1+g+g^2+\cdots )\\~\\ =-\frac{1}{2}-\frac{13 x}{2}-\frac{285 x^2}{8}+\mathcal{O}(x^3) \] where \(g = \frac{1}{4}(x^3+23x^2+20x) \)

ganeshie8 (ganeshie8):

I am satisfied with above work... but i would like to know if there is any other pleasant way to expand...

OpenStudy (trojanpoem):

@mathmale , The professor was meant to solve this e.g ( + other e.g) with us. Unluckily the lecture was canceled.

OpenStudy (trojanpoem):

@ganeshie8 , I am sure we will do some simplification before getting on. see this: sqrt(1+sinx) -> I solved it like :using the common expansion for (1+x)^m than plugging x = sinx Got it, but long solution. any shorter one ?

ganeshie8 (ganeshie8):

Thats pretty much how I would do it.. Have you also plugged in the series of sinx ?

OpenStudy (trojanpoem):

Yeah :(( and simplified

ganeshie8 (ganeshie8):

looks good, i can't think of anything better...

OpenStudy (trojanpoem):

I also thought of sqrt ( 1 + cos(pi/2 - x)) = sqrt(1 +2cos^2(pi/4 - x/2) - 1) = sqrt(2) cos(pi/4 - x/2) cos(pi/4 - x/2) isn't mac :/